A golf ball is hit from 4.3 m above a golfing fairway with an initial velocity of 30.0 m/s at an angle of 35° above the horizontal.
1
Expert's answer
2020-11-17T11:05:45-0500

"Solution"

"Given:\ y_D=4.3m\\
V_i=30m\ \ /s\ and\ \theta=35\degree\\
Where=V_{ix}=V_i\ cos\theta=(30m/s)\times\ Cos\ (35)\degree=24.57m/s\\
V_(iy)=V_i\ sin\theta=30m/s\times\ sin(35\degree)=17.207m/s"

i) In case we want to determine the time of flight of the ball then this can be solved as below;

Using the equation

"y=y_i+V_{iy}t+\frac{1}{2}a_yt^2\\
\therefore\ 0=4.3+(17.207)t+\frac{1}{2}(-9.81)t^2\\
(4.905)t^2-(17.207)t-4.3=0\\
Solving\ the\ quadratic\ equation\ for\ t\ gives\ t=3.742\ sec\\

ii) Also,\ the\ range\ of\ the\ golf\ ball\ can\ be\ determined\ as\ follows\\
\Delta\ d_x=V_{ix}\times\ t=(24.57m/s)(3.742sec)\\
\therefore\ \Delta\ d_x=91.94m\\"


iii) We can equally determine the velocity for the golf ball the instant before the ball impacts the ground by;

"Here,\ V_{fx}=V_{ix=24.57m/s}\\
while,\ V_{iy}+a_yt=(17.207)+(-9.81)(3.742)=-19.5m/s\\
Therefore\ V_f=\sqrt(24.57)^2+(-19.5)^2\\
=31.37m/s"